# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v3

@@ Conclusion
We presented Multi-Preference Lambda-weighted Listwise DPO, a generalization of Direct Preference Optimization (DPO) that enables alignment of language models with diverse and dynamic human preferences. By extending DPO from a pairwise to a listwise formulation and incorporating multiple preference dimensions via a lambda-weighted aggregation, our approach captures richer supervision signals and allows for controllable alignment without retraining.
Unlike traditional DPO, which assumes a single binary comparison per training instance, Lambda-DPO models the full distribution over ranked outputs, supporting scenarios where user intent is nuanced, multi-objective, or even conflicting. This enables practical applications such as personalized assistants that can adapt to individual user trade-offs (e.g., prioritizing helpfulness over conciseness), regulatory-sensitive deployments that balance safety and informativeness, and modular alignment pipelines where preferences may evolve over time.
Theoretically, our method is grounded in a KL-regularized RL objective, and empirically, we demonstrate competitive or superior performance to standard DPO across alignment benchmarks. By reducing estimation variance and providing interpretable preference control via the lambda vector, Lambda-DPO offers a scalable and expressive solution for real-world preference alignment.
In future work, we plan to explore more structured lambda scheduling, integration with curriculum learning, and extensions to conversational or multimodal alignment tasks.
@@ Acknowledgements
We would like to thank the contributors of the original DPO framework for their foundational work, which inspired this extension. Special thanks to the Hugging Face team and the LLaMA-Factory project for providing robust tools and open-source infrastructure for preference-based fine-tuning. We also acknowledge valuable discussions with colleagues and reviewers who provided feedback during the early stages of this work. Their insights have significantly improved the clarity and rigor of our theoretical analysis.
This work was supported in part by academic research funding and compute resources generously provided by [Your Institution Name Here].
@@ Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
The function thus reduces the four-way UltraFeedback preferences to a single pairwise comparison per prompt and optimizes the model via DPO.

### v4

@@ Conclusion
We introduced Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to a listwise setting with controllable $\lambda$ vectors for multi-dimensional alignment. This approach enables flexible, interpretable optimization over preference trade-offs without retraining.
Experiments on 1B–2B scale models show consistent gains over standard DPO, with average accuracy improvements of +1.7% on LLaMA 3.2-1B Instruct and +2.8% on Qwen-1.8 Chat. We further propose a learned scheduler that adaptively samples high-performing $\lambda$ configurations, outperforming uniform baselines while reducing modeling bias. These results highlight the effectiveness of listwise, preference-aware alignment in instruction tuning—especially for efficiently improving smaller models where preference supervision has greater impact, capacity is limited, and retraining is impractical. This is particularly relevant for domain-specific assistants on edge devices, multilingual student models distilled from larger systems, or rapid adaptation via methods like LoRA. In such settings, dynamically adjusting alignment objectives—e.g., helpfulness vs. honesty—without retraining offers both efficiency and controllability.
@@ Acknowledgements
We thank the contributors of the original DPO framework for their foundational work, which inspired this extension. We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao. We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
@@ Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
E.5 Unimodal Ablation
We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency. Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
Figure 9: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension. All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average. This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.